OpenStudy (anonymous):

Find the GCF of the terms of the polynomial. 20x4y + 25x2 A. 100x4y B. 5x2 C. 4x2 D. 5x4y Can someone explain how to do this problem? I got D but I don't think its right.

jimthompson5910 (jim_thompson5910):

5 is the GCF of 20 and 25, so you are on the right track.

jimthompson5910 (jim_thompson5910):

But the problem is with the variable term x^4y

jimthompson5910 (jim_thompson5910):

x^4 = x*x*x*x and x^2 = x*x In the terms above, how many times does x show up?

OpenStudy (anonymous):

2 times..? I'm not sure whether your talking about the one the 4 or the 2 or all together.

OpenStudy (anonymous):

it'll help you !!

jimthompson5910 (jim_thompson5910):

well in x^4, there are 4 copies of x and in x^2 there are 2 copies of x

jimthompson5910 (jim_thompson5910):

in both, there are two copies of x. This is the most copies of x they both share together

jimthompson5910 (jim_thompson5910):

that's why the GCF of x^4 and x^2 is x^2

jimthompson5910 (jim_thompson5910):

as for the y variables, there is a y in the first term but not in the second

jimthompson5910 (jim_thompson5910):

so y is NOT part of the GCF

OpenStudy (anonymous):

ohh now I get it. At first I was confused with the Y and the X but now I understand it :) Thank you!

jimthompson5910 (jim_thompson5910):

so what is the overall GCF?

OpenStudy (anonymous):

5x2 ?

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

Thank you so much! :)

jimthompson5910 (jim_thompson5910):

btw, you type 5x^2 to mean \(\Large 5x^2\), but I know what you meant to say

jimthompson5910 (jim_thompson5910):

you're welcome